• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Adaq

    Log entry time 01:59:55 on December20,2003

    Entry number 114653

    keyword=magnet woes

    Was contacted by software on-call (D. Wetherholt) about problem with Hall A
    momentum calculation not working properly. Arrived on-site and discovered that
    constants in momentum calculation for 6 degree septum were incorrect. With the
    help of J. Gomez, I fixed this problem and a couple of other associated
    problems with constants. After these modifications, the Q2 and Q3 power
    supplies ramped correctly to the momentum setpoint. The dipole and Q1 magnets
    still did not work properly. These problems appear to be hardware problems.
    The dipole uses the NMR to ramp to a desired setpoint. The NMR is unable to
    lock to a value. It appears that there is a problem with it's compensating
    coil. The Q1 appears to have a bit that is stuck. The Q1 magnet setpoint of
    1216 gives a readback of 1209, 1214 gives a readback of 1207, 1212 gives a
    readback of 1205, etc. The Q1 magnet does not give the correct readback until
    a setpoint of 1208 is requested. 1218 also gives a correct readback. J. Gomez
    informed the Hall A run coordinator of these problems. Hall A staff will
    address them.